### v1

@@ B.1 Model Selection - Per Language Results
For completeness, we present the individual benchmark results of the 1B-parameter model trained on 119B tokens for each language in the following tables: Table 9 for Chinese, Table 10 for French, Table 11 for German, Table 12 for Arabic, and Table 13 for Danish.
@@
B.2 Threshold Selection
To confirm that the CS filtering method is not competitive with MLP and FT, even when a higher percentage of documents is retained, we present the complete threshold selection results, including the CS method, in Table 14 in addition to the results shown in Section 4.2.2 (Table 2).
We provide further results on the variation in the average length of documents retained by our model-based filtering approaches for Chinese, French, Arabic, and Danish. These results complement the findings for German discussed in Section 4.2.2 and are shown in Figure 4. Table 15 lists the actual dataset sizes (number of retained tokens) after tokenization for all languages.
@@ B.2 Threshold Selection
| Baseline | 1597B | 730B | 973B | 127B | 108B |
B.3 Training Data Analysis
@@
B.4 Impact on multilingual model training
@@
Appendix C List of evaluation benchmarks and metrics
We provide a detailed overview of the evaluation benchmarks used to assess our models’ performance, along with their respective evaluation metrics in Table 20. For non-English tasks and English MMLU, we use the cloze multiple-choice prompt, which allows the model to directly predict each option instead of using the standard prompt format with A/B/C/D letter prefixes as targets. This approach was chosen because it has been shown to serve as a more reliable performance indicator earlier in training (Kydlíček et al., 2024). We evaluate the models in a 0-shot setting.
@@ Appendix C List of evaluation benchmarks and metrics
| AGIEval (Zhong et al., 2023) | ✓ | Normalized accuracy | |||||
| AlGhafa ARC (Almazrouei et al., 2023) | ✓ | Normalized accuracy | |||||
| AlGhafa PIQA (Almazrouei et al., 2023) | ✓ | Normalized accuracy | |||||
| AlGhafa RACE (Almazrouei et al., 2023) | ✓ | Normalized accuracy | |||||
| AlGhafa SciQ (Almazrouei et al., 2023) | ✓ | Normalized accuracy | |||||
| ARC (Clark et al., 2018) | ✓ | Normalized accuracy | |||||
| ARCD (Mozannar et al., 2019) | ✓ | F1 score | |||||
| Belebele (Bandarkar et al., 2024) | ✓ | ✓ | ✓ | ✓ | ✓ | Normalized accuracy | |
| C3 (Sun et al., 2020) | ✓ | Normalized accuracy | |||||
| C-Eval (Huang et al., 2023) | ✓ | Normalized accuracy | |||||
| Chinese-SQuAD (Pluto-Junzeng, 2019) | ✓ | F1 score | |||||
| CMMLU (Li et al., 2024a) | ✓ | Normalized accuracy | |||||
| CMRC 2018 (Cui et al., 2019) | ✓ | F1 score | |||||
| CommonsenseQA (Talmor et al., 2019) | ✓ | Normalized accuracy | |||||
| EXAMS (Hardalov et al., 2020) | ✓ | Normalized accuracy | |||||
| FQuAD (d’Hoffschmidt et al., 2020) | ✓ | F1 score | |||||
| HellaSwag (Zellers et al., 2019) | ✓ | Normalized accuracy | |||||
| M3Exam (Zhang et al., 2023) | ✓ | Normalized accuracy | |||||
| Mintaka (Sen et al., 2022) | ✓ | ✓ | F1 score | ||||
| MLMM ARC (Lai et al., 2023) | ✓ | ✓ | ✓ | Normalized accuracy | |||
| MLMM HellaSwag (Lai et al., 2023) | ✓ | ✓ | ✓ | ✓ | ✓ | Normalized accuracy | |
| MLMM MMLU (Lai et al., 2023) | ✓ | ✓ | ✓ | Normalized accuracy | |||
| MLQA (Lewis et al., 2020) | ✓ | F1 score | |||||
| MMLU (Hendrycks et al., 2020) | ✓ | Normalized accuracy | |||||
| OCNLI (Hu et al., 2020) | ✓ | Normalized accuracy | |||||
| OpenBookQA (Mihaylov et al., 2018) | ✓ | Normalized accuracy | |||||
| PIQA (Bisk et al., 2019) | ✓ | Normalized accuracy | |||||
| SOQAL (Mozannar et al., 2019) | ✓ | Normalized accuracy | |||||
| TriviaQA (Joshi et al., 2017) | ✓ | Quasi-exact match | |||||
| TyDi QA (Clark et al., 2020) | ✓ | F1 score | |||||
| WinoGrande (Sakaguchi et al., 2019) | ✓ | Normalized accuracy | |||||
| X-CODAH (Lin et al., 2021a) | ✓ | ✓ | ✓ | ✓ | Normalized accuracy | ||
| XCOPA (Ponti et al., 2020) | ✓ | Normalized accuracy | |||||
| X-CSQA (Lin et al., 2021a) | ✓ | ✓ | ✓ | ✓ | Normalized accuracy | ||
| XNLI 2.0 (Upadhyay & Upadhya, 2023) | ✓ | ✓ | ✓ | Normalized accuracy | |||
| XStoryCloze (Lin et al., 2021b) | ✓ | ✓ | Normalized accuracy | ||||
| XWINO (Tikhonov & Ryabinin, 2021) | ✓ | Normalized accuracy |
Appendix D FineWeb documents in different scoring approaches
@@
D.1 FastText Classifier (FT)
D.2 Multi-Layer Perceptron (MLP)
D.3 Cosine Similarity (CS)
Appendix E Example of a contaminated document

### v2

@@ C.1 Model Selection - Per Language Results
For clarity, we present the individual benchmark results of the 1B-parameter model trained on 119B tokens for each language in the following tables: Table 9 for Chinese, Table 10 for French, Table 11 for German, Table 12 for Arabic, and Table 13 for Danish.
@@
C.2 Threshold Selection
Complete Result. To confirm that the CS filtering method is not competitive with MLP and FT, even when a higher percentage of documents is retained, we present the complete threshold selection results, including the CS method, in Table 14 in addition to the results shown in Section 4.2.2 (Table 2).
Document Length Bias. Motivated by the observed bias in certain approaches favoring the selection of shorter documents, as seen in Figure 3, Figure 4 and Table 15, we examine how this bias interacts with performance when retaining more documents. As demonstrated in Table 15, the MLP MKC approach shows a tendency to retain shorter documents, while achieving higher performance with an increased number of retained documents. In contrast, the CS and FT filtering methods present mixed results, suggesting that the optimal threshold selection may be influenced by additional factors.
@@ C.2 Threshold Selection
[…]
C.3 Training Data Analysis
@@
C.4 Replay of Original Data
We explore whether incorporating a small percentage of original raw data (replay) can help improve performance. We do this for our best FastText (FT MKC+) and Transformer approaches (MLP MKC+). Table 16 presents the results of experiments where 5% and 10% unfiltered data were mixed into the training dataset, alongside results from training without any replay. Although, the FT MKC+ filters shows mixed signal, our MLP MKC+ approach clearly demonstrates that replay does not improve performance, indicating the data selection already retains enough diversity. In cases of less diverse datasets, replay was shown to offer benefits [Bethune et al., 2025, Chen et al., 2023].
| Approach | Mixture Rate | Average Rank |
| MLP MKC+ | 5% | 5.09 |
[…]
| MLP MKC+ | 10% | 5.40 |
| FT MKC+ | 10% | 7.17 |
| FT MKC+ | 0% | 7.51 |
| FT MKC+ | 5% | 8.66 |
C.5 Impact on multilingual model training
@@
Appendix D List of evaluation benchmarks and metrics
We provide a detailed overview of the evaluation benchmarks used to assess our models’ performance, along with their respective evaluation metrics in Table 21. For non-English tasks and English MMLU, we use the cloze multiple-choice prompt, which allows the model to directly predict each option instead of using the standard prompt format with A/B/C/D letter prefixes as targets. This approach was chosen because it has been shown to serve as a more reliable performance indicator earlier in training [Kydlíček et al., 2024]. We evaluate the models in a 0-shot setting.
@@ Appendix D List of evaluation benchmarks and metrics
| AGIEval [Zhong et al., 2023] | ✓ | Normalized accuracy | |||||
| AlGhafa ARC [Almazrouei et al., 2023] | ✓ | Normalized accuracy | |||||
| AlGhafa PIQA [Almazrouei et al., 2023] | ✓ | Normalized accuracy | |||||
| AlGhafa RACE [Almazrouei et al., 2023] | ✓ | Normalized accuracy | |||||
| AlGhafa SciQ [Almazrouei et al., 2023] | ✓ | Normalized accuracy | |||||
| ArabicMMLU [Koto et al., 2024] | ✓ | Normalized accuracy | |||||
| ARC [Clark et al., 2018] | ✓ | Normalized accuracy | |||||
| ARCD [Mozannar et al., 2019] | ✓ | F1 score | |||||
| Belebele [Bandarkar et al., 2024] | ✓ | ✓ | ✓ | ✓ | ✓ | Normalized accuracy | |
| C3 [Sun et al., 2020] | ✓ | Normalized accuracy | |||||
| C-Eval [Huang et al., 2023] | ✓ | Normalized accuracy | |||||
| Chinese-SQuAD [Pluto-Junzeng, 2019] | ✓ | F1 score | |||||
| CMMLU [Li et al., 2024a] | ✓ | Normalized accuracy | |||||
| CMRC 2018 [Cui et al., 2019] | ✓ | F1 score | |||||
| CommonsenseQA [Talmor et al., 2019] | ✓ | Normalized accuracy | |||||
| EXAMS [Hardalov et al., 2020] | ✓ | Normalized accuracy | |||||
| FQuAD [d’Hoffschmidt et al., 2020] | ✓ | F1 score | |||||
| HellaSwag [Zellers et al., 2019] | ✓ | Normalized accuracy | |||||
| M3Exam [Zhang et al., 2023] | ✓ | Normalized accuracy | |||||
| Meta MMLU [Llama Team, 2024] | ✓ | ✓ | Normalized accuracy | ||||
| Mintaka [Sen et al., 2022] | ✓ | ✓ | F1 score | ||||
| MLMM ARC [Lai et al., 2023] | ✓ | ✓ | ✓ | Normalized accuracy | |||
| MLMM HellaSwag [Lai et al., 2023] | ✓ | ✓ | ✓ | ✓ | ✓ | Normalized accuracy | |
| MLQA [Lewis et al., 2020] | ✓ | F1 score | |||||
| MMLU [Hendrycks et al., 2020] | ✓ | Normalized accuracy | |||||
| OCNLI [Hu et al., 2020] | ✓ | Normalized accuracy | |||||
| OpenBookQA [Mihaylov et al., 2018] | ✓ | Normalized accuracy | |||||
| PIQA [Bisk et al., 2019] | ✓ | Normalized accuracy | |||||
| SOQAL [Mozannar et al., 2019] | ✓ | Normalized accuracy | |||||
| TriviaQA [Joshi et al., 2017] | ✓ | Quasi-exact match | |||||
| TyDi QA [Clark et al., 2020] | ✓ | F1 score | |||||
| WinoGrande [Sakaguchi et al., 2019] | ✓ | Normalized accuracy | |||||
| X-CODAH [Lin et al., 2021a] | ✓ | ✓ | ✓ | ✓ | Normalized accuracy | ||
| XCOPA [Ponti et al., 2020] | ✓ | Normalized accuracy | |||||
| X-CSQA [Lin et al., 2021a] | ✓ | ✓ | ✓ | ✓ | Normalized accuracy | ||
| XNLI 2.0 [Upadhyay and Upadhya, 2023] | ✓ | ✓ | ✓ | Normalized accuracy | |||
| XStoryCloze [Lin et al., 2021b] | ✓ | ✓ | Normalized accuracy | ||||
| XWINO [Tikhonov and Ryabinin, 2021] | ✓ | Normalized accuracy |
Appendix E Average Rank Computation
Analogous to the method in FineTasks [Kydlíček et al., 2024], we compute the average rank for our ablations as follows:
1.
We train a model for each parameter configuration we want to ablate on.
2.
We evaluate each model on all the selected benchmarks.
3.
We compute the rank of each model (individual experiment) with regard to each benchmark and language.
4.
We compute the average rank for each model across all benchmarks and languages.
Appendix F FineWeb documents in different scoring approaches
@@
F.1 FastText Classifier (FT)
F.2 Multi-Layer Perceptron (MLP)
F.3 Cosine Similarity (CS)
Appendix G Example of a contaminated document
@@
Appendix H License Information
H.1 Dataset Licenses
We use the following pretraining datasets:
FineWeb-2 (ODC-By license)
FineWeb (ODC-By license)
FineWeb-Edu (ODC-By license)
DCLM (CC-BY 4.0)
We use the following classifier training datasets:
Aya Collection (Apache 2.0 license)
Aya Dataset (Apache 2.0 license)
Translated multilingual MMLU [OpenAI, 2024] (MIT license)
OpenAssistant-2 (Apache 2.0 license)
Include-Base-44 (Apache 2.0 license)
H.2 Code Licenses
We use the following open source code:
Nanotron (Apache 2.0 license)
Datatrove (Apache 2.0 license)
Lighteval (MIT license)
FastText (MIT license)
H.3 Model Licenses
We use the following models:
Mistral v3 (Tekken) (Apache 2.0 license)
XLM-RoBERTa (MIT license)